# Fluctuation Theorem and Optimal Control of an Active Tracking Particle with Information Processing

### v3

@@ Abstract
Living systems often function with regulatory interactions, but the question of how activity, stochasticity and regulations work together for achieving different goals still remains puzzling. We propose the model of an active tracking particle with information processing, based on which the entropy production, information flow, and generalised fluctuation theorem are derived. Moreover, the system performance, in terms of the first passage steps and the total energy consumption, are analysed in the variable space of (measurement error, control field), leading to discussions on the optimal control of the system. Not only elucidating the basic concepts involved in a stochastic active system with information processing, this prototypical model could also inspire more elaborated modelings of natural smart organisms and industrial designs of controllable active systems with desired physical performances in the future.
@@
In this work, we propose the model of an active tracking particle with information processing, based on a Bayesian description. With the model, we obtain the generalised fluctuation theorem, serving as the basic obeying principle of the information flow and entropy production. Moreover, we investigate how the performance of the system, in terms of the first passage steps and the energy consumptions, can depend on the information-related processes including measurements and feedback controls, providing the optimal control protocols in the corresponding parameter space.
@@
For seeking the transparency of the model, we take a one-dimensional and descritised description to illustrate how a regulated active particle should move, as shown in Fig. 1. The state of the active particle at any time, $t_{k\in\mathbb{Z}}$, is described by two variables $(x_{k},n_{k})$, where $x_{k}=s~\!\Delta x$ denotes its location with $s\in\mathbb{Z}$ and $\Delta x$ as the step length, and $n_{k}\in(L,R)$ denotes its orientation pointing to the left or the right; without the information processing, the model reduces to the run-and-tumble model describing the dynamics of a bacterium [38, 39]. The regulatory purpose of the active particle is to move from the location $x=0$ to the destination $x=\mathcal{D}\Delta x$ along a straight line, i.e., the location should be $x=d~\!\Delta x$ at any time $t_{d}$, and this is a typical tracking problem in control theories. To achieve feedback controls, we consider the particle carrying a magnetic moment of magnitude $m$ in the same direction as its orientation, and the orientation can be controlled by applying an external magnetic field; such a model can be easily generalised to other controlling cases, e.g., by utilising light signal on algae [17, 18] or chemical concentration on bacteria [40, 41]. As shown in Fig. 1, the particle starts moving from state $(x_{0},n_{0})$ at time $t_{0}$, and after measurement - feedback control- active motion - relaxation, the particle reaches a new state $(x_{1},n_{1})$ at time $t_{1}$. By repeating the above process, the particle stops moving after arriving at the destination $(\mathcal{D}\Delta x,n_{\mathcal{T}})$ after $\mathcal{T}$ time steps ($\mathcal{T}\geq\mathcal{D}$); note that the particle motion can deviate from the desired trajectory (straight line) due to the intrinsic stochasticity of the system.
@@
We utilise a bipartite Bayesian network [42, 43] to describe the dynamics of the active particle, consisting of a physical subsystem $\mathcal{X}$ in charge of the particle locomotion as an ordinary run-and-tumble particle and a control (measurement - feedback) subsystem $\mathcal{C}$ in charge of the information-based regulation. With this treatment, the regulation of the whole process simply follows the classical Bellman description [44] of dynamic programming where the total optimal cost is the summation of all the costs at every time step, so it is sufficient to focus on the regulation process during a single time step, $t_{k}\rightarrow t_{k+1}$ with $0\leq k<\mathcal{T}$. For achieving regulatory purposes during any arbitrary time step, $t_{k}\rightarrow t_{k+1}$, we take the Maxwell-demon-like setup of the measurement - feedback process [45, 46, 47], and it is detailed as follows. At time $t_{k}$, the state of the particle is $(x_{k},n_{k})$, where $n_{k}=L,R$ with equal probability at equilibrium, i.e., $p_{\mathrm{eq}}^{0}(n_{k}=L,R)=1/2$. Then we measure its orientation denoted as $m_{k}$, and such measurements have an error, $0\leq\epsilon\leq 1$; in other words, we can define a conditional probability characterising the measurement: $p(m_{k}|n_{k})=1-\epsilon$ for $m_{k}=n_{k}$ and $p(m_{k}|n_{k})=\epsilon$ for $m_{k}\neq n_{k}$. If the measured orientation is pointing to the left, i.e., $m_{k}=L$, then we apply a magnetic field of the magnitude $B$ and the direction pointing to the right during a finite time interval $\Delta t$ to change its orientation; if the measured orientation is pointing to the right, i.e., $m_{k}=R$, then we do not apply the magnetic field and let the particle orientation freely evolve during $\Delta t$ by thermal fluctuation. We use $n_{k}^{\prime}$ to denote the particle orientation after the measurement and feedback control. The transition, $n_{k}\rightarrow n_{k}^{\prime}$, which is critical in the regulatory process, includes the nodes, $n_{k}$, $m_{k}$, and $n_{k}^{\prime}$ as shown in Fig. 1, and the transition can be described by the conditional probability, $p(n^{\prime}_{k}|n_{k},m_{k})$, with its explicit expressions shown in the table of the Appendix. For example, no magnetic field is applied in the case of ($n_{k}=L$, $m_{k}=R$), where the particle orientation will evolve freely during $\Delta t$, and we can obtain the probabilities of finding the particle with $n_{k}^{\prime}=L$ and $n_{k}^{\prime}=R$ as, $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$ and $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$, respectively, where the transition rate is $k_{0}=k_{0}^{L\rightarrow R}+k_{0}^{R\rightarrow L}=2k_{0}^{L\rightarrow R}$ with $k_{0}^{L\rightarrow R}$ and $k_{0}^{R\rightarrow L}$ as the transition rate from $n_{k}=L$ to $n_{k}^{\prime}=R$ and that from $n_{k}=R$ to $n_{k}^{\prime}=L$ without magnetic field, respectively. In following discussions, the notation $u_{p}=\exp[-k_{0}\Delta t]$ (related with unchanged population) is taken for simplicity. After the measurement - feedback process, we let the particle move along its regulated orientation with a constant step length, $\Delta x$, which is described by the process $(x_{k},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k}^{\prime})$, and such dynamic process can be achieved in various active systems, e.g., by shining light on photocatalytical colloidal particles [48, 49] or applying electric field on metal-dielectric Janus particles [50, 51]. Then let the particle orientation relax to the state with its orientation distribution recovering to $p_{\mathrm{eq}}^{0}(n_{k+1}=L,R)=1/2$ at time $t_{k+1}$. For those familiar with bacteria swimming in the run-and-tumble mode, the above setup can be treated as a bacterium moving with a ‘brain’ which performs measurement and feedback control. By repeating the above process from $t_{0}$ to $t_{\mathcal{T}}$, the particle can arrive at its destination after $\mathcal{T}$ time steps, with its state as $(\mathcal{D}\Delta x,n_{\mathcal{T}})$.
Entropy production, information flow and fluctuation theorem.
The total entropy production in the whole system consists of the entropy production in both the subsystem $\mathcal{X}$ and its connected heat bath [42, 52], i.e., $\Delta\mathcal{S}_{\mathrm{tot}}=\Delta\mathcal{S}_{\mathcal{X}}+\Delta\mathcal{S}_{\mathrm{bath}}=\sum_{k=0}^{\mathcal{T}-1}(\Delta s_{\mathcal{X}}^{k}+\Delta s_{\mathrm{bath}}^{k})$, where $\Delta s_{\mathcal{X}}^{k}$ and $\Delta s_{\mathrm{bath}}^{k}$ denote the entropy production of the subsystem $\mathcal{X}$ and the heat bath during the time step, $t_{k}\rightarrow t_{k+1}$, respectively. During any time step, $t_{k}\rightarrow t_{k+1}$, the entropy production of the subsystem $\mathcal{X}$ is simply,
$\Delta s_{\mathcal{X}}^{k}=\ln\frac{p(n_{k})}{p(n_{k+1})},$ (1)
where $p(n_{k})$ denotes the probability distribution function of the particle orientation at time $t_{k}$. For explicit expressions of relevant probabilities in the system, one can refer to supplemental materials [53]. Regarding the entropy change of the heat bath, $\Delta s_{\mathrm{bath}}^{k}$, during the time step, $t_{k}\rightarrow t_{k+1}$, which is due to the heat dissipated into the heat bath from the system $\mathcal{X}$, there are two parts: one during the process $(x_{k},n_{k})\rightarrow(x_{k},n_{k}^{\prime})$, and the other during the process $(x_{k+1},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k+1})$, which are denoted by $\Delta s_{\mathrm{bath}}^{k1}$ and $\Delta s_{\mathrm{bath}}^{k2}$, respectively. The entropy production, $\Delta s_{\mathrm{bath}}^{k1}$, during the process, $(x_{k},n_{k})\rightarrow(x_{k},n_{k}^{\prime})$, can be calculated as:
$\Delta s_{\mathrm{bath}}^{k1}=-\beta Q_{k}^{1}=\ln\frac{p(n^{\prime}_{k}|n_{k},m_{k})}{p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})},$ (2)
where $\beta=1/k_{B}T$ with $T$ as the temperature, $Q_{k}^{1}$ denotes the dissipated heat during the process $(x_{k},n_{k})\rightarrow(x_{k},n_{k}^{\prime})$, and $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})$ denotes the backward transition probability from $\hat{n}^{\prime}_{k}$ to $\hat{n}_{k}$ depending on the memory $\hat{m}_{k}$, with $\hat{n}_{k}$, $\hat{n}^{\prime}_{k}$ and $\hat{m}_{k}$ as time-reversed $n_{k}$, $n^{\prime}_{k}$ and $m_{k}$, respectively. The backward transition probability, $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})$, can be related with the forward transition probability, $p(n^{\prime}_{k}|n_{k};m_{k})$, generally as: $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k};\lambda_{k})/p(n^{\prime}_{k}|n_{k};\lambda_{k})=\exp[-\beta E(n_{k};\lambda_{k})+\beta E(n^{\prime}_{k};\lambda_{k})]$, where $\lambda_{k}$ is external control parameter characterising the control on $\mathcal{X}$ and $E(n_{k};\lambda_{k})$ denotes the system energy. Without magnetic field, we have degenerated states: $E(n_{k};\lambda_{k})=E(n^{\prime}_{k};\lambda_{k})=0$, and there are relations, $\exp[-\beta E(n_{k};\lambda_{k})]=\exp[-\beta E(n^{\prime}_{k};\lambda_{k})]=1$ and $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k};\lambda_{k})=p(n^{\prime}_{k}|n_{k};\lambda_{k})$. With magnetic field, the degeneracy is lifted, and we have $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k};\lambda_{k})\neq p(n^{\prime}_{k}|n_{k};\lambda_{k})$, whose values depend on $\hat{n}_{k}$ and $\hat{n}^{\prime}_{k}$. The entropy production, $\Delta s_{\mathrm{bath}}^{k2}$, during another process, $(x_{k+1},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k+1})$, is simply:
$\Delta s_{\mathrm{bath}}^{k2}=-\beta Q_{k}^{2}=\ln\frac{p(n_{k+1}|n^{\prime}_{k})}{p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})}.$ (3)
Then the total entropy production during the time step, $t_{k}\rightarrow t_{k+1}$, becomes:
$\sigma^{k}$ $=$ $\Delta s_{\mathcal{X}}^{k}+\Delta s_{\mathrm{bath}}^{k1}+\Delta s_{\mathrm{bath}}^{k2}$ (4)
$=$ $\ln\frac{p(n_{k})p(n^{\prime}_{k}|n_{k},m_{k})p(n_{k+1}|n^{\prime}_{k})}{p(n_{k+1})p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})}.$
One can easily prove that, if there is no measurement and feedback control, the total entropy production satisfies the integral fluctuation theorem: $\langle e^{-\Delta\mathcal{S}_{\mathrm{tot}}}\rangle=\langle e^{-\sum_{k=0}^{\mathcal{T}-1}\sigma^{k}}\rangle=\prod_{k=0}^{\mathcal{T}-1}\left[\sum_{\hat{n}_{k},\hat{n}^{\prime}_{k},n_{k+1}}p(n_{k+1})p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k})p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})\right]=1$, where $\langle$$\rangle$ indicates the ensemble average and the last equality is guaranteed by the normalisation condition. Then there is the second law of thermodynamics: $\langle\Delta\mathcal{S}_{\mathrm{tot}}\rangle\geq 0$ obtained from Jensen’s inequality. By considering the measurement and feedback control, the relation $\langle\Delta\mathcal{S}_{\mathrm{tot}}\rangle\geq 0$ may not hold any more. For characterising how measurement and feedback control can influence the dynamics of the active particle, we introduce an information quantity, $\Theta_{k}$ [42]:
$\Theta_{k}=I_{k}^{\mathrm{fin}}-I_{k}^{\mathrm{tr}}-I_{k}^{\mathrm{ini}},$ (5)
@@
where the mutual information $I_{k}^{\mathrm{fin}}=I(n^{\prime}_{k}:m_{k})=\ln[p(n^{\prime}_{k},m_{k})/p(n^{\prime}_{k})p(m_{k})]$ characterises the correlation between $n^{\prime}_{k}$ and $m_{k}$, the mutual information $I_{k}^{\mathrm{tr}}=I(m_{k}:n_{k})=\ln[p(m_{k},n_{k})/p(m_{k})p(n_{k})]$ denotes the transfer entropy from the physical system $\mathcal{X}$ to the control system $\mathcal{C}$, and $I_{k}^{\mathrm{ini}}=0$ denotes no correlation between $n_{k}$ and previous measurements, $m_{l<k}$. By summing them up, we obtain the dynamic information flow in the system, $\Theta_{d}=\sum_{k=0}^{\mathcal{T}-1}\Theta_{k}$. One can prove that the total entropy production, $\Delta\mathcal{S}_{\mathrm{tot}}$, and the dynamic information flow, $\Theta_{d}$, satisfy the generalised integral fluctuation theorem:
$\langle e^{-\Delta\mathcal{S}_{\mathrm{tot}}+\Theta_{d}}\rangle=$ (6)
$~\!\!\prod_{k=0}^{\mathcal{T}-1}\Big{[}\!\!\!\sum_{n_{k},m_{k}}^{n^{\prime}_{k},n_{k+1}}\frac{p_{B}(\hat{n}_{k+1},\hat{n}^{\prime}_{k},\hat{m}_{k},\hat{n}_{k})}{p(n_{k},m_{k},n^{\prime}_{k},n_{k+1})}p(n_{k},m_{k},n^{\prime}_{k},n_{k+1})\!\Big{]}\!=\!1,$
as obtained in other information-incorporated stochastic systems [54, 55, 56, 42], where the forward probability is $p(n_{k},m_{k},n^{\prime}_{k},n_{k+1})\quad=\quad p(n_{k})p(m_{k}|n_{k})p(n^{\prime}_{k}|n_{k},m_{k})\times\\ p(n_{k+1}|n^{\prime}_{k})$ and the backward probability is $p_{B}(\hat{n}_{k},\hat{m}_{k},\hat{n}^{\prime}_{k},\hat{n}_{k+1})=p(n_{k+1})p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})p(m_{k}|n^{\prime}_{k})\times\\ p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})$. Alternatively, one can utilise the Kullback-Leibler (KL) divergence, $d_{\mathrm{KL}}=\Delta\mathcal{S}_{\mathrm{tot}}-\Theta_{d}=\ln[p(\mathcal{X}\bigcup\mathcal{C})/p_{B}(\mathcal{X}\bigcup\mathcal{C})]$, to express the generalised fluctuation theorem: $\langle e^{-d_{\mathrm{KL}}}\rangle$=1. By again applying Jensen’s inequality, we can obtain the generalised second law of thermodynamics, $\langle\Delta\mathcal{S}_{\mathrm{tot}}\rangle\geq\langle\Theta_{d}\rangle$ [42, 57], indicating that the total entropy production $\langle\Delta\mathcal{S}_{\mathrm{tot}}\rangle$ can be negative.
@@
Figure 3: Mean energy consumption per time step, $\ln[\langle\mathcal{E}\rangle/(\mathcal{D}E_{A}^{0})]$, as a function of the measurement error and the control field for parameters of (a) $\mathcal{B}=1/2,\mathcal{M}=1$, (b) $\mathcal{B}=1/2$, $\mathcal{M}=30$, and (c) $\mathcal{B}=10$, $\mathcal{M}=0.1$, where the black circles denote the local minima and the red triangle denote the global minimum. (d) Optimal measurement error and (e) optimal field as a function of the measurement parameter, $\mathcal{M}$, for the control parameter of $\mathcal{B}=1/2$ (black line), 1010 (red line) and 5050 (blue line). Squares and circles in (e) denote the values of the optimal field at the limit of $\mathcal{M}\rightarrow 0$ and $\mathcal{M}\rightarrow\infty$, respectively. (f) Optimal field at the limit of $\mathcal{M}\rightarrow 0~(\epsilon=0)$, $\hat{B}^{\mathrm{opt}}_{\epsilon=0}$, and (g) at the limit of $\mathcal{M}\rightarrow\infty~(\epsilon=1)$, $\hat{B}^{\mathrm{opt}}_{\epsilon=1}$, as a function of the control parameter, $\mathcal{B}$. Inset of (f) denote the values of the optimal field for the full range of $\mathcal{B}\in(10^{-6},5\times 10^{1})$.
@@
which depends on the process of measurement and feedback control through the probability, $p(n^{\prime}_{k}=L)$ [explicit expression shown in the table of the Appendix]. Note that $\langle\mathcal{T}\rangle$ diverges for unbiased random walk with $p(n^{\prime}_{k}=R)=1/2$. Apparently, the averaged first passage steps decreases with increasing magnetic field and decreasing measurement error, as shown in Fig. 2. We introduce the dimensionless field strength, $\hat{B}=B/B_{0}$ with $B_{0}=k_{B}T/m$ as the referenced magnetic field. For weak field controls, $\hat{B}\rightarrow 0$, the averaged first passage steps decreases with the increasing magnetic field in the form of $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/[\hat{B}(1-u_{p})]$, and for strong field controls, $\hat{B}\rightarrow\infty$, the averaged first passage steps only depends on the measurement error, in the form of $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/(1+u_{p}-2\epsilon u_{p})$.

### v4

# arXiv:2508.21487v4[cond-mat.soft] 27 Feb 2026

@@ Abstract
Living systems often function with regulatory interactions, but the question of how activity, stochasticity and regulations work together for achieving different goals still remains puzzling. We propose a stochastic model of an active tracking particle with information processing, where the entropy production and information flow are discussed, with the generalised fluctuation theorem serving as a benchmark for verifying the probability setups. Based on the model, the system performance, in terms of the first passage steps and the total energy consumption, are analysed in the variable space of (measurement error, control field), leading to discussions on optimal controls of the system. Not only elucidating the basic concepts involved in a stochastic active system with information processing, this prototypical model could also inspire more elaborated modelings of natural smart organisms and industrial designs of controllable active systems with desired physical performances in the future.
@@
In this work, we propose a stochastic model of an active tracking particle with information processing, based on the Bayesian description. We discuss about the entropy production and the information flow of the system, where the generalised fluctuation theorem serves as a benchmark for verifying the probability setups. With the model, we investigate how the system performance, in terms of the first passage steps and the total energy consumption, can depend on the information-related processes including measurements and feedback controls, providing the optimal control protocols in the corresponding parameter space.
@@
For seeking the transparency of the model, we take a one-dimensional and descritised description to illustrate how a regulated active particle should move, as shown in Fig. 1. The state of the active particle at any time, $t_{k\in\mathbb{Z}}$, is described by two variables $(x_{k},n_{k})$, where $x_{k}=s~\!\Delta x$ denotes its location with $s\in\mathbb{Z}$ and $\Delta x$ as the step length, and $n_{k}\in(L,R)$ denotes its orientation pointing to the left or the right; without the information processing, the model reduces to the run-and-tumble model describing the dynamics of a bacterium [38, 39]. The regulatory purpose of the active particle is to move from the location $x=0$ to the destination $x=\mathcal{D}\Delta x$ along a straight line without any step moving to the left, i.e., the location should be $x=d~\!\Delta x$ at any time $t_{d}$, and this is a typical tracking problem in control theories. To achieve feedback controls, we consider the particle carrying a magnetic moment of magnitude $m$ in the same direction as its orientation, and the orientation can be controlled by applying an external magnetic field; such a model can be easily generalised to other controlling cases, e.g., by utilising light signal on algae [17, 18] or chemical concentration on bacteria [40, 41]. As shown in Fig. 1, the particle starts moving from state $(x_{0},n_{0})$ at time $t_{0}$, and after measurement - feedback control- active motion - relaxation, the particle reaches a new state $(x_{1},n_{1})$ at time $t_{1}$. By repeating the above processes, the particle stops moving after arriving at the destination $(\mathcal{D}\Delta x,n_{\mathcal{T}})$ after $\mathcal{T}$ time steps ($\mathcal{T}\geq\mathcal{D}$); note that the particle motion can deviate from the desired trajectory (keep moving to the right without any step moving to the left) due to the intrinsic stochasticity of the system.
@@
We utilise a bipartite Bayesian network [42, 43] to describe the dynamics of the active particle, consisting of a physical subsystem $\mathcal{X}$ in charge of the particle locomotion as an ordinary run-and-tumble particle and a control (measurement - feedback) subsystem $\mathcal{C}$ in charge of the information-based regulation. Note that the system ($\mathcal{X}$ + $\mathcal{C}$) is placed in a heat bath, subjected to thermal fluctuations. With this treatment, the regulation of the whole process simply follows the classical Bellman description [44] of dynamic programming where the total optimal cost is the summation of all the costs at every time step, so it is sufficient to focus on the regulation process during a single time step, $t_{k}\rightarrow t_{k+1}$ with $0\leq k<\mathcal{T}$. For achieving regulatory purpose during any arbitrary time step, $t_{k}\rightarrow t_{k+1}$, we take the Maxwell-demon-like setup of the measurement - feedback process [45, 46, 47], and it is detailed as follows. At time $t_{k}$, the state of the particle is $(x_{k},n_{k})$, where $n_{k}=L,R$ with equal probability at equilibrium, i.e., $p(n_{k}=L,R)=p_{\mathrm{eq}}^{0}(n_{k}=L,R)=1/2$. Then we measure its orientation denoted as $m_{k}$, and such measurements have an error (probability for a wrong measurement) [48, 49, 50], $0\leq\epsilon\leq 1$, which may arise from limited instrument precision, thermal fluctuation, etc.; in other words, we can define a conditional probability characterising this measurement: $p(m_{k}|n_{k})=1-\epsilon$ for $m_{k}=n_{k}$ and $p(m_{k}|n_{k})=\epsilon$ for $m_{k}\neq n_{k}$. If the measured orientation is pointing to the left, i.e., $m_{k}=L$, then we apply a magnetic field of the magnitude $B$ and the direction pointing to the right during a finite time interval $\Delta t$ to change its orientation; if the measured orientation is pointing to the right, i.e., $m_{k}=R$, then we do not apply the magnetic field and let the particle orientation freely evolve during $\Delta t$ driven by thermal fluctuation. We use $n_{k}^{\prime}$ to denote the particle orientation after the measurement and feedback control. The transition, $n_{k}\rightarrow n_{k}^{\prime}$, which is critical in the regulatory process, includes the nodes, $n_{k}$, $m_{k}$, and $n_{k}^{\prime}$ as shown in Fig. 1, and the transition can be described by the conditional probability, $p(n^{\prime}_{k}|n_{k},m_{k})$, with its explicit expressions shown in Table 1 of the End Matter, by mapping this dynamic process to a typical two-state Kramers process [51, 52, 53]. For example, no magnetic field is applied in the case of ($n_{k}=L$, $m_{k}=R$), where the particle orientation will evolve freely during $\Delta t$, and we can obtain the probabilities of finding the particle with $n_{k}^{\prime}=L$ and $n_{k}^{\prime}=R$ as, $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$ and $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$, respectively, where the transition rate is $k_{0}=k_{0}^{L\rightarrow R}+k_{0}^{R\rightarrow L}=2k_{0}^{L\rightarrow R}$ with $k_{0}^{L\rightarrow R}$ and $k_{0}^{R\rightarrow L}$ as the transition rate from $n_{k}=L$ to $n_{k}^{\prime}=R$ and that from $n_{k}=R$ to $n_{k}^{\prime}=L$ without magnetic field, respectively. In following discussions, the notation $u_{p}=\exp[-k_{0}\Delta t]$ (related with unchanged population) is taken for simplicity. When applying magnetic field, the probability, $p(n^{\prime}_{k}|n_{k},m_{k})$, depends explicitly on the strength of the magnetic field due to thermal fluctuations, as shown Table 1. After the measurement - feedback process, we let the particle move along its regulated orientation with a constant step length, $\Delta x$, which is described by the process $(x_{k},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k}^{\prime})$, i.e., for the ‘move’ process (analogous to the ‘run’ step of a run-and-tumble bacterium) in Fig. 1(b), the spatial position of the particle will become $x_{k+1}=x_{k}+\Delta x$ for $n^{\prime}_{k}=R$, and become $x_{k+1}=x_{k}-\Delta x$ for $n^{\prime}_{k}=L$. Such dynamic process can be achieved in various active systems, e.g., by applying magnetic field on platinum-nickel-gold nanorods [54], shining light on gold-polystyrene Janus particles [55] or applying electric field on metal-dielectric Janus particles [56, 57]. After this, we let the particle orientation relax to the equilibrium state with its orientation distribution as $p_{\mathrm{eq}}^{0}(n_{k+1}=L,R)=1/2$ at time $t_{k+1}$. For those familiar with bacteria swimming in the run-and-tumble mode, the above setup can be treated as a bacterium moving with a ‘brain’ which performs measurement and feedback control. By repeating the above processes from $t_{0}$ to $t_{\mathcal{T}}$, the particle can arrive at its destination after $\mathcal{T}$ time steps, with its state as $(\mathcal{D}\Delta x,n_{\mathcal{T}})$.
Entropy production and information flow.
The total entropy production in the whole system consists of the entropy production in both the subsystem $\mathcal{X}$ and its connected heat bath [42, 58], i.e., $\Delta\mathcal{S}_{\mathrm{tot}}=\Delta\mathcal{S}_{\mathcal{X}}+\Delta\mathcal{S}_{\mathrm{bath}}=\sum_{k=0}^{\mathcal{T}-1}(\Delta s_{\mathcal{X}}^{k}+\Delta s_{\mathrm{bath}}^{k})$, where $\Delta s_{\mathcal{X}}^{k}$ and $\Delta s_{\mathrm{bath}}^{k}$ denote the entropy production of the subsystem $\mathcal{X}$ and the heat bath during the time step, $t_{k}\rightarrow t_{k+1}$, respectively. The entropy production of the subsystem $\mathcal{X}$, $\Delta s_{\mathcal{X}}^{k}=\ln[p(n_{k})/p(n_{k+1})]=0$, since particle orientation at time $t_{k}$ and $t_{k+1}$ are both in the equilibrium state, i.e., $p(n_{k})=p(n_{k+1})=1/2$. The entropy change of the heat bath, $\Delta s_{\mathrm{bath}}^{k}$, includes two contributions, $\Delta s_{\mathrm{bath}}^{k1}$ during the ‘evolution’ process of $(x_{k},n_{k})\rightarrow(x_{k},n_{k}^{\prime})$, and $\Delta s_{\mathrm{bath}}^{k2}$ during the ‘relax’ process of $(x_{k+1},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k+1})$. The first contribution is $\Delta s_{\mathrm{bath}}^{k1}=\ln[p(n^{\prime}_{k}|n_{k},m_{k})/p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})]$, where $p(n^{\prime}_{k}|n_{k},m_{k})$ denotes the transitional probability as introduced above, and $p_{B}(\hat{n}_{k}|\hat{n}^{\prime}_{k},\hat{m}_{k})$ denotes the backward transition probability from $\hat{n}^{\prime}_{k}$ to $\hat{n}_{k}$ depending on the memory $\hat{m}_{k}$, with $\hat{n}_{k}$, $\hat{n}^{\prime}_{k}$ and $\hat{m}_{k}$ as time-reversed $n_{k}$, $n^{\prime}_{k}$ and $m_{k}$, respectively. The second contribution is $\Delta s_{\mathrm{bath}}^{k2}=\ln[p(n_{k+1}|n^{\prime}_{k})/p_{B}(\hat{n}^{\prime}_{k}|\hat{n}_{k+1})]$, denoting the relaxation process of the particle orientation, and it is always positive. For explicit expressions of the above mentioned probabilities, one can refer to the supplemental materials for more details [51]. Note that the entropy production during the ‘move’ process is zero, i.e., the ‘move’ process is reversible by reversing the direction of the particle, which is different from the case of a particle driven by a constant force without changing the driving direction [51].
@@
For characterising how measurement and feedback control can influence the dynamics of the active particle, we apply the information quantity, $\Theta_{k}=I_{k}^{\mathrm{fin}}-I_{k}^{\mathrm{tr}}-I_{k}^{\mathrm{ini}}$ [42], where the mutual information $I_{k}^{\mathrm{fin}}=I(n^{\prime}_{k}:m_{k})=\ln[p(n^{\prime}_{k},m_{k})/p(n^{\prime}_{k})p(m_{k})]$ characterises the correlation between $n^{\prime}_{k}$ and $m_{k}$, the mutual information $I_{k}^{\mathrm{tr}}=I(m_{k}:n_{k})=\ln[p(m_{k},n_{k})/p(m_{k})p(n_{k})]$ denotes the transfer entropy from the physical system $\mathcal{X}$ to the control system $\mathcal{C}$, and $I_{k}^{\mathrm{ini}}=0$ denotes no correlation between $n_{k}$ and previous measurements, $m_{l<k}$. By summing them up, we obtain the dynamic information flow in the system, $\Theta_{d}=\sum_{k=0}^{\mathcal{T}-1}\Theta_{k}$. One can prove that the total entropy production, $\Delta\mathcal{S}_{\mathrm{tot}}$, and the dynamic information flow, $\Theta_{d}$, satisfy the generalised integral fluctuation theorem: $\langle e^{-\Delta\mathcal{S}_{\mathrm{tot}}+\Theta_{d}}\rangle=1$, as obtained in other information-incorporated stochastic systems [59, 60, 61, 62, 63, 42, 64], which is utilised as a benchmark in this work for verifying the probability setup of the above dynamic processes.
@@
Figure 3: Mean energy consumption per time step, $\ln[\langle\mathcal{E}\rangle/(\mathcal{D}E_{A}^{0})]$, as a function of the measurement error and the control field for the parameters of (a) $\mathcal{B}=1/2,\mathcal{M}=1$, (b) $\mathcal{B}=1/2$, $\mathcal{M}=30$, and (c) $\mathcal{B}=10$, $\mathcal{M}=0.1$, where the black circles denote the local minima and the red triangle denote the global minimum. The control parameter, $\mathcal{B}=E_{B}^{0}/E_{A}^{0}$, compares the magnetic-control energy scale to the active (running) energy $E_{A}^{0}$, and the measurement parameter, $\mathcal{M}=E_{M}^{0}/E_{A}^{0}$, compares the measurement energy scale to the active energy $E_{A}^{0}$. (d) Optimal measurement error and (e) optimal field as a function of the measurement parameter, $\mathcal{M}$, for the control parameter of $\mathcal{B}=1/2$ (black line), 1010 (red line) and 5050 (blue line). Squares and circles in (e) denote the values of the optimal field at the limit of $\mathcal{M}\rightarrow 0$ and $\mathcal{M}\rightarrow\infty$, respectively. (f) Optimal field at the limit of $\mathcal{M}\rightarrow 0~(\epsilon=0)$, $\hat{B}^{\mathrm{opt}}_{\epsilon=0}$, and (g) at the limit of $\mathcal{M}\rightarrow\infty~(\epsilon=1)$, $\hat{B}^{\mathrm{opt}}_{\epsilon=1}$, as a function of the control parameter, $\mathcal{B}$. Inset of (f) denote the values of the optimal field for the full range of $\mathcal{B}\in(10^{-6},5\times 10^{1})$.
@@
which depends on the process of measurement and feedback control through the probability, $p(n^{\prime}_{k}=R)$ [explicit expressions shown in the Table 2 of the End Matter]. Note that $\langle\mathcal{T}\rangle$ diverges for unbiased random walk with $p(n^{\prime}_{k}=R)=1/2$. Apparently, the averaged first passage steps decreases with increasing magnetic field and decreasing measurement error, as shown in Fig. 2. We introduce the dimensionless field strength, $\hat{B}=B/B_{0}$ with $B_{0}=k_{B}T/m$ as the referenced magnetic field. For weak field controls, $\hat{B}\rightarrow 0$, the averaged first passage steps decreases with the increasing magnetic field in the form of $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/[\hat{B}(1-u_{p})]$, and for strong field controls, $\hat{B}\rightarrow\infty$, the averaged first passage steps only depends on the measurement error, in the form of $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/(1+u_{p}-2\epsilon u_{p})$.